MC Qu. 5-A1 What is the profitability index for an investment with the follow...

What is the profitability index for an investment with the following cash flows given a 8 percent required return?

Year         Cash Flow
0 $-21,000   
1 $7,100   
2 $9,400   
3 $8,900   
Answer #1

Let Cash Flows in Year n be denoted by CFn
Given,
CF0 = -21000
CF1 = 7100
CF2 = 9400
CF3 = 8900
Interest Rate = r = 0.08
Net Present Value = PV = ΣCFn/(1+r)n
= -21000 + 7100/(1+0.08) + 9400/(1+0.08)2 + 8900/(1+0.08)3
= $698.17

Profitability Index = Present Value of Future Cash Flows / Initial Investment

= (NPV + Initial Investment) / Initial Investment

= (698.17 + 21000)/21000

= 1.033